# A Theoretical Framework for Grokking: Interpolation followed by Riemannian Norm Minimisation

### v1

@@ 1 Introduction
Strikingly simple algorithms such as gradient methods are a driving force behind the success of deep learning. Nonetheless, their remarkable performance remains mysterious, and a full theoretical understanding is lacking. In particular: (i) convergence to low training loss solutions on non-convex objectives is far from trivial, and (ii) it is unclear why the resulting solutions generalise well (Zhang et al., 2017). These questions are accompanied by a range of surprising phenomena that arise during training. One such intriguing behaviour is known as the grokking phenomenon, which we explore in this work. Coined by Power et al. (2022), this term describes a two-phase pattern in the learning curves: first, the training loss rapidly decreases to zero, while the test loss plateaus at a certain value. This is followed by a second phase, where the training loss remains zero, but the test loss steadily decreases, leading to a final improved generalisation performance, as depicted in Figure 1 (left).
In this paper, we propose a novel theoretical perspective to explain this phenomenon. By examining the gradient flow dynamics with weight decay, we show that, in the limit of vanishing weight decay, we can fully describe the trajectory of the model parameters. Specifically, we prove that the training process can be decomposed into two distinct phases. In the first phase, the gradient flow follows the unregularised path, converging to a manifold of critical points of the training loss. In the second, the trajectory enters a slow drift phase, where the weights move along this manifold, driven by weight decay, gradually reducing their $\ell_{2}$ norm, as illustrated in Figure 1 (right). We argue that this slow decrease in the weight norms explains the grokking phenomenon, as smaller weight norms are often correlated with better generalisation. To convey the main intuition, we state below an informal version of our result, describing the full trajectory of the gradient flow with small weight decay.
@@
As the weight decay parameter $\lambda$ is taken to 00, the trajectory $w^{\lambda}(t)$ can be seen as a composition of two coupled dynamics:
@@
    (Fast dynamics driven by F given by Proposition 1) In a first phase, the weights follow the unregularised gradient flow and converge to a manifold of critical points of $F$.
@@
    (Slow dynamics driven by the weight decay given by Proposition 2) At time $t\approx 1/\lambda$, the iterates start slowly drifting along this manifold, following a Riemannian gradient flow that decreases the $\ell_{2}$-norm of the weights.
@@
Note this is a purely optimisation result: no statistical assumptions are made, and it a priori does not imply any improvement in test loss during the slow phase. However, it provides a natural explanation for the grokking phenomenon. Indeed, in practice, for many deep learning models with random initialisation, gradient flow converges to a global minimiser of the training loss. When this solution generalises poorly—as is often the case with large initial weights, in the so-called lazy regime (Chizat et al., 2019)—the subsequent slow drift along the critical manifold, driven by weight decay, decreases the $\ell_{2}$-norm of the solution and simplifies it in the second phase. Since lower weight norms often correlate with better generalisation (Bach, 2017; Liu et al., 2022c; D’Angelo et al., 2024), this offers a convincing explanation for the delayed improvement in test performance. We discuss various settings where this behaviour is observed in Section 5.
Figure 1: Gradient flow with small weight decay $\lambda$. (Left) A typical example of grokking: the training loss rapidly drops to zero, while the test loss plateaus for a long period before eventually decreasing—coinciding with a steady drop in the $\ell_{2}$-norm of the weights. (Right) Schematic illustration in parameter space $\mathbb{R}^{d}$ of the optimisation behaviour described in Theorem 1. The trajectory $w^{\lambda}(t)$ initially follows the unregularised gradient flow and converges to a manifold of critical points of $F$ (fast dynamics). At time $t\approx 1/\lambda$, the regularisation term becomes dominant and induces a slow drift along this manifold toward a lower $\ell_{2}$-norm solution (slow dynamics).
@@ 2 Related work
The term grokking was originally coined by Power et al. (2022), which studied a two-layer transformer trained with weight decay on a modular addition task. They observed that the network quickly fits the training data while generalising poorly, followed much later by a sudden transition to near-perfect generalisation. Following this work, many studies have investigated modular addition tasks to better understand the mechanisms underlying this phenomenon (Nanda et al., 2023; Gromov, 2023). However, grokking has been observed far beyond this setting. For instance, Barak et al. (2022) showed that training a neural network to learn parities exhibits a similar delayed generalisation pattern. In Liu et al. (2022c), grokking was induced across a broad range of tasks, including image classification and sentiment analysis, by using small datasets, large initialisations, and weight decay. Other settings and architectures where grokking-like behaviour appears include matrix factorisation (Lyu et al., 2023) and learning XOR-clustered data with a ReLU network (Xu et al., 2023). Finally, it is worth noting that this delayed transition in generalisation was already observed in earlier works, as clearly illustrated in Figure 3 of Chizat and Bach (2020).
@@
Several works have framed grokking as the transition between the lazy and rich regimes. The lazy regime, also called the NTK regime, was introduced by Jacot et al. (2018). It typically arises when the network is trained from large initialisations (Chizat et al., 2019), and corresponds to a setting where zero training loss can be quickly achieved, but often with poor generalisation performance. In contrast, the rich regime (also called the feature learning regime) corresponds to a setting where the network actively learns new internal features during training. In the classification setting, Lyu and Li (2019) show that the rich regime is always attained for homogeneous parameterisations, and similarly, Chizat and Bach (2020) provide an analogous result for infinitely wide two-layer networks. In this context, Lyu et al. (2023) and Kumar et al. (2024), followed by Mohamadi et al. (2024), offer a theoretical perspective on grokking as the transition from the lazy regime to the rich regime during training: initially, the predictor quickly converges towards the NTK solution, and later escapes this regime to reach a better generalising solution, driven by the effects of implicit regularisation and/or weight decay.
@@
The role of weight decay in grokking remains somewhat debated. While many of the original works exhibiting the phenomenon include weight decay (Power et al., 2022; Liu et al., 2022b), grokking has also been observed without (Chizat and Bach, 2020; Xu et al., 2023), as strongly emphasised by Kumar et al. (2024). However, as shown in Lyu et al. (2023), the transition tends to occur much later and to be less sharp without weight decay. In this context, weight decay can be interpreted as a factor that triggers or accelerates the transition from the lazy to the rich regime. While grokking can be observed in classification tasks even without weight decay—thanks to the algorithm’s implicit bias—to the best of our knowledge, it cannot occur in regression tasks unless weight decay is used. Of particular relevance to our work, Liu et al. (2022c) propose an intuitive explanation of grokking that is based on weight decay: during the first phase, the model rapidly converges to a poor global minimum; during the second, slower phase, weight decay gradually steers the iterates toward a lower-norm solution with better generalisation properties. While appealing, this explanation remains informal and lacks rigorous theoretical support. In this work, we provide a formal analysis of the optimisation dynamics underlying the grokking phenomenon: an initial fast phase leads to convergence toward the solution associated with the lazy regime, followed by a slower second phase that drives convergence toward the solution characteristic of the rich regime.
@@
Many theoretical works in the machine learning community have studied the training dynamics of gradient methods in overparameterised neural networks, where the set of zero training loss solutions forms a high-dimensional manifold. In this context, leveraging results from dynamical systems theory (Katzenberger, 1990), the work of Li et al. (2021) describes the drift dynamics induced by stochastic noise after stochastic gradient descent (SGD) reaches the manifold. This analysis was further extended by Shalova et al. (2024). Leveraging a similar stochastic differential framework, Vivien et al. (2022) precisely characterises this drift in the setting of diagonal linear networks, and proves that it leads to desirable sparsity guarantees. Much in the spirit of our work, although outside the deep learning context, Fatkullin et al. (2010) derive stochastic differential equations that describe the dynamics of systems with small random perturbations on energy landscapes with manifolds of minima, illustrating how the system first converges to the manifold and then drifts along it. Whereas these works emphasise drift induced by the stochasticity of gradient updates, our analysis focuses on a deterministic drift arising from the regularisation of the training objective.
@@ 3 Setting and preliminaries
We consider a loss function $F:\mathbb{R}^{d}\to\mathbb{R}_{+}$ which is sufficiently smooth, as stated in Assumption 1 below. Typical examples include the least square loss over some training dataset, where the parameters to optimise represent the weights of some neural network architecture. For a given $\lambda>0$, we define the regularised loss $F_{\lambda}$ as:
@@
The regularity conditions on $F$ ensure that, for any $\lambda\geq 0$, the gradient flow ODE has a unique solution, which is defined for all $t\geq 0$. This is a consequence of the Picard–Lindelöf theorem and boundedness of the trajectories. Definability in the o-minimal sense guarantees that bounded gradient flow trajectories converge to a limit point (Kurdyka, 1998, Thm. 2). This is a mild assumption satisfied by most functions arising in applications, such as polynomials, logarithms, exponentials, subanalytic functions, and finite combinations of those; see Coste (1999); Bolte et al. (2007) for more details. Finally, note that the boundedness assumption on the unregularised flow excludes classification settings where the network can perfectly separate the data, since in such cases the unregularised iterates diverge.
@@ 3.1 Stationary manifold and Riemannian flow
Assumption 1 guarantees that the gradient flow converges to a limit point $w^{\mathrm{GF}}_{\infty}\coloneqq\lim_{t\to\infty}w^{\mathrm{GF}}(t)$, which is a stationary point of $F$. In typical scenarios of training overparameterised models, stationary points are not isolated, but form continuous sets (Cooper, 2021).
@@
Our key assumption is that $\mathcal{M}$ forms a smooth manifold, and that it contains only local minimizers (and not saddle points). Additionnally, we impose that the non-zero eigenvalues of the Hessian on $\mathcal{M}$ are lower bounded by some constant $\eta>0$. Following the terminology of Rebjock and Boumal (2024), this is known as the Morse-Bott property.
@@
As stated in Rebjock and Boumal (2024), for $\mathcal{C}^{2}$ functions this property is equivalent to the Kurdyka–Łojasiewicz, also known as Polyak–Łojasiewicz (PL) condition locally around $\mathcal{M}$ (Kurdyka, 1998; Bolte et al., 2009). It implies that (i) every point in $\mathcal{M}$ is a local minimiser, and (ii) all gradient flow trajectories are locally attracted towards $\mathcal{M}$. This stability property is essential for proving regularity of the flow map $\Phi$ in Section 4. Let us discuss the relevance of Assumption 2 in the context of overparameterised machine learning.
  • Convergence to a local minimiser: our assumption rules out the possibility that $w^{\mathrm{GF}}$ converges to a saddle point of $F$. This is justified by a large number of works showing that gradient methods avoid saddle points for almost all initialisations. In particular, Lee et al. (2016, 2019) prove it under the assumption that all saddle points of $F$ are strict. Although their result only holds for discrete-time gradient descent, the underlying argument can be extended to gradient flow (the proof relies on the Stable Manifold Theorem for dynamical systems, which also holds in continuous time (Teschl, 2012, §7)).
  • Morse-Bott/Łojasiewicz property: this is a common assumption in the analysis of gradient flow dynamics for overparameterised networks (Li et al., 2021; Fatkullin et al., 2010; Shalova et al., 2024). Note that for general models, the critical set may not form a manifold everywhere. However, it is often possible to show that the manifold structure holds on most of the space, excluding some degenerate points; see Section 5 for examples and Liu et al. (2022a) for a generic result. Moreover, the results derived from such an assumption are generally very representative of empirical observations, as can be seen in Section 5.
@@
where $P_{T_{\mathcal{M}}(w)}$ is the orthogonal projection on the tangent space to $\mathcal{M}$ at $w$. Under Assumption 2, typical properties of smooth manifolds imply that $T_{\mathcal{M}}(w)=\mathrm{Ker}(\nabla^{2}F(w))$ (see e.g., Boumal, 2023, for a detailed introduction to optimisation on manifolds). Using this notion of Riemannian gradient, we study the Riemannian gradient flow for some objective function $h$ and initialization $w_{\mathcal{M}}\in\mathcal{M}$, defined as the curve $w$ satisfying,
@@
By construction of the Riemannian gradient, the trajectory of any solution of this ODE necessarily belongs to $\nabla F^{-1}(0)$, and therefore to $\mathcal{M}$, since $\mathcal{M}$ is a maximal connected component. If $h$ is $\mathcal{C}^{2}$ and has compact sublevel sets, Assumptions 1 and 2 guarantee that there exists a unique solution to Equation 3 and that it is defined on $\mathbb{R}_{+}$.

### v2

@@ 1 Introduction
Strikingly simple algorithms such as gradient methods are a driving force behind the success of deep learning. Nonetheless, their remarkable performance remains mysterious, and a full theoretical understanding is lacking. In particular: (i) convergence to low training loss solutions on non-convex objectives is far from trivial, and (ii) it is unclear why the resulting solutions generalise well (recht2017understanding). These questions are accompanied by a range of surprising phenomena that arise during training. One such intriguing behaviour is known as the grokking phenomenon, which we explore in this work. Coined by power2022grokking, this term describes a two-phase pattern in the learning curves: first, the training loss rapidly decreases to zero, while the test loss plateaus at a certain value. This is followed by a second phase, where the training loss remains zero, but the test loss steadily decreases, leading to a final improved generalisation performance, as depicted in Figure˜1 (left).
In this paper, we propose a novel theoretical perspective to explain this phenomenon. By examining the gradient flow dynamics with weight decay, we show that, in the limit of vanishing weight decay, we can fully describe the trajectory of the model parameters. Specifically, we prove that the training process can be decomposed into two distinct phases. In the first phase, the gradient flow follows the unregularised path, converging to a manifold of critical points of the training loss. In the second, the trajectory enters a slow drift phase, where the weights move along this manifold, driven by weight decay, gradually reducing their $\ell_{2}$ norm, as illustrated in Figure˜1 (right). We argue that this slow decrease in the weight norms explains the grokking phenomenon, as smaller weight norms are often correlated with better generalisation. To convey the main intuition, we state below an informal version of our result, describing the full trajectory of the gradient flow with small weight decay.
@@
As the weight decay parameter $\lambda$ is taken to 0, the trajectory $w^{\lambda}(t)$ can be seen as a composition of two coupled dynamics:
@@
    (Fast dynamics driven by F given by Proposition˜1) In a first phase, the weights follow the unregularised gradient flow and converge to a manifold of critical points of $F$.
@@
    (Slow dynamics driven by the weight decay given by Proposition˜2) At time $t\approx 1/\lambda$, the iterates start slowly drifting along this manifold, following a Riemannian gradient flow that decreases the $\ell_{2}$-norm of the weights.
@@
Note this is a purely optimisation result: no statistical assumptions are made, and it a priori does not imply any improvement in test loss during the slow phase. However, it provides a natural explanation for the grokking phenomenon. Indeed, in practice, for many deep learning models with random initialisation, gradient flow converges to a global minimiser of the training loss. When this solution generalises poorly—as is often the case with large initial weights, in the so-called lazy regime (chizat2019lazy)—the subsequent slow drift along the critical manifold, driven by weight decay, decreases the $\ell_{2}$-norm of the solution and simplifies it in the second phase. Since lower weight norms often correlate with better generalisation (bach2017breaking; liu2022omnigrok; d2024we), this offers a convincing explanation for the delayed improvement in test performance. We discuss various settings where this behaviour is observed in Section˜5.
Figure 1: Gradient flow with small weight decay $\lambda$. (Left) A typical example of grokking: the training loss rapidly drops to zero, while the test loss plateaus for a long period before eventually decreasing—coinciding with a steady drop in the $\ell_{2}$-norm of the weights. (Right) Schematic illustration in parameter space $\mathbb{R}^{d}$ of the optimisation behaviour described in Theorem˜1. The trajectory $w^{\lambda}(t)$ initially follows the unregularised gradient flow and converges to a manifold of critical points of $F$ (fast dynamics). At time $t\approx 1/\lambda$, the regularisation term becomes dominant and induces a slow drift along this manifold toward a lower $\ell_{2}$-norm solution (slow dynamics).
@@ 2 Related work
The term grokking was originally coined by power2022grokking, which studied a two-layer transformer trained with weight decay on a modular addition task. They observed that the network quickly fits the training data while generalising poorly, followed much later by a sudden transition to near-perfect generalisation. Following this work, many studies have investigated modular addition tasks to better understand the mechanisms underlying this phenomenon (nanda2023progress; gromov2023grokking). However, grokking has been observed far beyond this setting. For instance, barak2022hidden showed that training a neural network to learn parities exhibits a similar delayed generalisation pattern. In liu2022omnigrok, grokking was induced across a broad range of tasks, including image classification and sentiment analysis, by using small datasets, large initialisations, and weight decay. Other settings and architectures where grokking-like behaviour appears include matrix factorisation (lyu2023dichotomy) and learning XOR-clustered data with a ReLU network (xu2023benign). Finally, it is worth noting that this delayed transition in generalisation was already observed in earlier works, as clearly illustrated in Figure 3 of chizat2020implicit. More recently, jeffares2025position argued that grokking may not be so central to Deep Learning and may only appear in very specific situations. However, we still believe its a priori counter-intuitive aspect is worth investigating and might lead to theoretical understandings that go beyond what is currently referred by grokking.
@@
Several works have framed grokking as the transition between the lazy and rich regimes. The lazy regime, also called the NTK regime, was introduced by jacot2018neural. It typically arises when the network is trained from large initialisations (chizat2019lazy), and corresponds to a setting where zero training loss can be quickly achieved, but often with poor generalisation performance. In contrast, the rich regime (also called the feature learning regime) corresponds to a setting where the network actively learns new internal features during training. In the classification setting, lyu2019gradient show that the rich regime is always attained for homogeneous parameterisations, and similarly, chizat2020implicit provide an analogous result for infinitely wide two-layer networks. In this context, lyu2023dichotomy and kumar2024grokking, followed by mohamadi2024you, offer a theoretical perspective on grokking as the transition from the lazy regime to the rich regime during training: initially, the predictor quickly converges towards the NTK solution, and later escapes this regime to reach a better generalising solution, driven by the effects of implicit regularisation and/or weight decay.
@@
The role of weight decay in grokking remains somewhat debated. While many of the original works exhibiting the phenomenon include weight decay (power2022grokking; liu2022towards), grokking has also been observed without (chizat2020implicit; xu2023benign), as strongly emphasised by kumar2024grokking. However, as shown in lyu2023dichotomy, the transition tends to occur much later and to be less sharp without weight decay. In this context, weight decay can be interpreted as a factor that triggers or accelerates the transition from the lazy to the rich regime. While grokking can be observed in classification tasks even without weight decay—thanks to the algorithm’s implicit bias—to the best of our knowledge, it cannot occur in regression tasks unless weight decay is used. Of particular relevance to our work, liu2022omnigrok propose an intuitive explanation of grokking that is based on weight decay: during the first phase, the model rapidly converges to a poor global minimum; during the second, slower phase, weight decay gradually steers the iterates toward a lower-norm solution with better generalisation properties. While appealing, this explanation remains informal and lacks rigorous theoretical support. In this work, we provide a formal analysis of the optimisation dynamics underlying the grokking phenomenon: an initial fast phase leads to convergence toward the solution associated with the lazy regime, followed by a slower second phase that drives convergence toward the solution characteristic of the rich regime.
@@
Many theoretical works in the machine learning community have studied the training dynamics of gradient methods in overparameterised neural networks, where the set of zero training loss solutions forms a high-dimensional manifold. In this context, leveraging results from dynamical systems theory (katzenberger1990solutions), the work of li2021happens describes the drift dynamics induced by stochastic noise after stochastic gradient descent (SGD) reaches the manifold. This analysis was further extended by shalova2024singular. Leveraging a similar stochastic differential framework, vivien2022label precisely characterises this drift in the setting of diagonal linear networks, and proves that it leads to desirable sparsity guarantees. Much in the spirit of our work, although outside the deep learning context, fatkullin2010reduced derive stochastic differential equations that describe the dynamics of systems with small random perturbations on energy landscapes with manifolds of minima, illustrating how the system first converges to the manifold and then drifts along it.
Our analysis builds upon this framework popularized by li2021happens and tracing back to katzenberger1990solutions. Our contribution differs from previous work in both focus and scope. Whereas previous analyses attribute the second, slower learning phase to stochastic effects, we identify a deterministic mechanism–namely, a slow drift induced by regularization–and link it directly to the grokking phenomenon, a connection that, to our knowledge, has not been previously explored. From a technical perspective, our setting is simpler yet enables a more detailed analysis. Rather than appealing to results from katzenberger1990solutions as a black box, we provide a simpler, self-contained proof building on falconer1983differentiation. Furthermore, unlike prior analyses that assume initialization near the interpolation manifold, our framework accommodates arbitrary initialisation. We rigorously characterise the initial convergence toward the manifold and the subsequent transition to the drift phase along it, which constitutes the main technical novelty of our analysis.
@@ 3 Setting and preliminaries
We consider a loss function $F:\mathbb{R}^{d}\to\mathbb{R}_{+}$ which is sufficiently smooth, as stated in Assumption˜1 below. Typical examples include the least square loss over some training dataset, where the parameters to optimise represent the weights of some neural network architecture. For a given $\lambda>0$, we define the regularised loss $F_{\lambda}$ as:
@@
The regularity conditions on $F$ ensure that, for any $\lambda\geq 0$, the gradient flow ODE has a unique solution, which is defined for all $t\geq 0$. This is a consequence of the Picard–Lindelöf theorem and boundedness of the trajectories. Definability in the o-minimal sense guarantees that bounded gradient flow trajectories converge to a limit point (kurdyka1998gradient, Thm. 2). This is a mild assumption satisfied by most functions arising in applications, such as polynomials, logarithms, exponentials, subanalytic functions, and finite combinations of those; see coste2000introduction; Bolte2007 for more details. Overall, Assumption˜1 is pretty mild and holds for all architectures (e.g., neural networks or transformers) that use differentiable activations.
Finally, note that the boundedness assumption on the unregularised flow excludes classification settings where the network can perfectly separate the data, since in such cases the unregularised iterates diverge.
@@ 3.1 Stationary manifold and Riemannian flow
Assumption˜1 guarantees that the gradient flow converges to a limit point $w^{\mathrm{GF}}_{\infty}\coloneqq\lim_{t\to\infty}w^{\mathrm{GF}}(t)$, which is a stationary point of $F$. In typical scenarios of training overparameterised models, stationary points are not isolated, but form continuous sets (Cooper2021global).
@@
Our key assumption is that $\mathcal{M}$ forms a smooth manifold, and that it contains only local minimizers (and not saddle points). Additionnally, we impose that the non-zero eigenvalues of the Hessian on $\mathcal{M}$ are lower bounded by some constant $\eta>0$. Following the terminology of rebjock2024fast, this is known as the Morse-Bott property.
@@
As stated in rebjock2024fast, for $\mathcal{C}^{2}$ functions this property is equivalent to the Kurdyka–Łojasiewicz, also known as Polyak–Łojasiewicz (PL) condition locally around $\mathcal{M}$ (kurdyka1998gradient; Bolte2009). It implies that (i) every point in $\mathcal{M}$ is a local minimiser, and (ii) all gradient flow trajectories are locally attracted towards $\mathcal{M}$. This stability property is essential for proving regularity of the flow map $\Phi$ in Section˜4. Let us discuss the relevance of Assumption˜2 in the context of overparameterised machine learning.
  • Convergence to a local minimiser: our assumption rules out the possibility that $w^{\mathrm{GF}}$ converges to a saddle point of $F$. This is justified by a large number of works showing that gradient methods avoid saddle points for almost all initialisations. In particular, Jordan2016saddleavoidance; Lee2019 prove it under the assumption that all saddle points of $F$ are strict. Although their result only holds for discrete-time gradient descent, the underlying argument can be extended to gradient flow (the proof relies on the Stable Manifold Theorem for dynamical systems, which also holds in continuous time (teschl2012ordinary, §7)).
  • Morse-Bott/Łojasiewicz property: this is a common assumption in the analysis of gradient flow dynamics for overparameterised networks (li2021happens; fatkullin2010reduced; shalova2024singular). Note that for general models, the critical set may not form a manifold everywhere. However, it is often possible to show that the manifold structure holds locally, i.e., on most of the space, excluding some degenerate points [^1] While our results are stated for such a global Morse-Bott assumption, they could be directly extended to local Morse-Bott assumptions, provided that the iterates remain in the non-degenerate region.; see Section˜5 for examples and liu2022loss for a generic result. Moreover, the results derived from such an assumption are generally very representative of empirical observations, as can be seen in Section˜5.
@@
where $P_{T_{\mathcal{M}}(w)}$ is the orthogonal projection on the tangent space to $\mathcal{M}$ at $w$. Under Assumption˜2, typical properties of smooth manifolds imply that $T_{\mathcal{M}}(w)=\mathrm{Ker}(\nabla^{2}F(w))$ (see e.g., boumal2023introduction, for a detailed introduction to optimisation on manifolds). Using this notion of Riemannian gradient, we study the Riemannian gradient flow for some objective function $h$ and initialization $w_{\mathcal{M}}\in\mathcal{M}$, defined as the curve $w$ satisfying,
@@
By construction of the Riemannian gradient, the trajectory of any solution of this ODE necessarily belongs to $\nabla F^{-1}(0)$, and therefore to $\mathcal{M}$, since $\mathcal{M}$ is a maximal connected component. If $h$ is $\mathcal{C}^{2}$ and has compact sublevel sets, Assumptions 1 and 2 guarantee that there exists a unique solution to Equation˜3 and that it is defined on $\mathbb{R}_{+}$.